OpenStudy (anonymous):

Solve. 2x^2 + 3x = 20.

OpenStudy (anonymous):

x = –10 and x = 3 x = five over two and x= –4 x = 5 and x = 2 x = –five over two and x = –4

OpenStudy (kc_kennylau):

2x^2+3x-20=0 (2x-5)(x+4)=0 2x-5=0 or x+4=0 x=0.4 or x=-4

OpenStudy (anonymous):

thank u

OpenStudy (kc_kennylau):

no problem :)